OpenStudy (anonymous):

For y = x2 + 6x − 16, Determine if the parabola opens up or down. State if the vertex will be a maximum or minimum. Find the vertex. Find the x-intercepts. Describe the graph of the equation.

OpenStudy (anonymous):

leading coefficient is positive ( it is 1) so it opens up

OpenStudy (anonymous):

that makes the second coordinate of the vertex a minimum

OpenStudy (anonymous):

well, ahah yeah that part was easy, i should have left that our. @satellite73

OpenStudy (anonymous):

find the vertex by knowing that the first coordinate is \(-\frac{b}{2a}\) which in your case is \[-\frac{6}{2\times 1}=-3\]

OpenStudy (anonymous):

 ..?